### v1

@@ 8 Example of Emergency Braking Control
Figure 20 and 21 show state trajectories and feasible regions of RL under HJ reachability constraint at different iterations. As is the case in other constraints, the EFR monotonically expands to the maximum one, and the state trajectories converge to the optimal ones. Like safety index, the IFR under HJ reachability constraint is much larger than the EFR at an early stage of training, which is also because of the minimum restrictiveness on the next state.
Figure 20: State trajectories of RL under HJR constraint at different iterations.
Figure 21: Feasible regions of RL under HJR constraint at different iterations.
9 Conclusion
@@
This paper proposes a feasibility theory that applies to both MPC and RL. Compared with existing theories built for MPC, our theory fills in the missing part of feasibility analysis for an arbitrary policy. Based on a decoupling view of virtual-time domain and real-time domain, we separately define initial and endless, state and policy feasibility, and their corresponding feasible regions. We further analyze the containment relationships between different feasible regions, which enables us to analyze the feasibility of an arbitrary policy. After that, we provide virtual-time constraint design rules along with a practical design tool called feasibility function that helps achieve the maximum feasible region. The feasibility function is categorized into control invariant set and constraint aggregation, which lead to different kinds of virtual-time constraints. We review most of existing constraint formulations and point out that they are essentially applications of feasibility functions in different forms. Finally, we demonstrate our feasibility theory in an emergency braking control task by visualizing initially and endlessly feasible regions of MPC and RL policies under different kinds of virtual-time constraints.

### v2

@@ 8.1 Emergency braking control
Figure 8.15 and 8.16 show state trajectories and feasible regions of RL under HJ reachability constraint at different iterations. As is the case in other constraints, the EFR monotonically expands to the maximum one, and the state trajectories converge to the optimal ones. Like SI, the IFR under HJ reachability constraint is much larger than the EFR at an early stage of training, which is also because of the minimum restrictiveness on the next state.
Figure 8.15: State trajectories of RL under HJR constraint at different iterations.
Figure 8.16: Feasible regions of RL under HJR constraint at different iterations.
8.2 Unicycle obstacle avoidance
In this task, a unicycle aims to move forward while avoiding a circular obstacle in the 2D plane, as shown in Figure 8.17.
Figure 8.17: Unicycle obstacle avoidance scenario.
The unicycle has a four-dimensional state $x_{t}=[y_{t},z_{t},v_{t},\theta_{t}]^{\top}$ and a two-dimensional action $u_{t}=[a_{t},\omega_{t}]^{\top}$, where $(y,z)$ is the coordinate on the 2D plane, $v$ is the velocity, $\theta$ is the heading angle, $a$ is the acceleration, and $\omega$ is the angular velocity. The system dynamics are given as follows:
$y_{t+1}$ $=y_{t}+\Delta t\cdot v_{t}\cos\theta_{t}$ (8.16)
$z_{t+1}$ $=z_{t}+\Delta t\cdot v_{t}\sin\theta_{t}$
$v_{t+1}$ $=v_{t}+\Delta t\cdot a_{t}$
$\theta_{t+1}$ $=\theta_{t}+\Delta t\cdot\omega_{t}$
The time step is $\Delta t=0.1\mathrm{s}$. The actions are bounded as $a\in[-1,1]\ \mathrm{m/s^{2}}$ and $\omega\in[-\pi/4,\pi/4]\ \mathrm{rad/s}$. The reward function is defined as
$r(x_{t},u_{t})=z_{t},$ (8.17)
which encourages forward progress in the $z$-direction. The constraint is defined to avoid collision with the obstacle:
$h(x_{t})=r-\left\lVert[y_{t},z_{t}]^{\top}-p_{\mathrm{obs}}\right\rVert_{2}\leq 0,$ (8.18)
where $r=0.5\mathrm{m}$ and $p_{\mathrm{obs}}=[0,0]^{\top}$ is the diameter and the center of the obstacle, respectively.
We consider three types of virtual-time constraints: (1) pointwise constraint, (2) CBF constraint, and (3) HJ reachability constraint. The pointwise constraint follows its standard formulation. The CBF is designed as
$B(x_{i|t})=0.7-d_{i|t}-k\cdot\dot{d}_{i|t},$ (8.19)
where $k>0$ is a tunable parameter, and $d_{i|t}$ is the distance between the unicycle and the obstacle, with its rate of change computed as
$\dot{d}_{i|t}=v_{i|t}\cos(\theta_{i|t}-\arctan(z_{i|t}/y_{i|t})).$ (8.20)
The HJ reachability function cannot be computed analytically for this system. Instead, we use a neural network to approximate it and jointly train it with an RL policy using the feasible policy iteration [yang2023feasible] framework. Training the HJ reachability function relies on the risky self-consistenty condition introduced in Section 6.3.2. We follow the practice of fisac2019bridging, which introduces a discount factor $\gamma\in(0,1)$ to the self-consistency condition to construct a contraction mapping:
$F^{\pi}(x)=(1-\gamma)h(x)+\gamma\max\{h(x),F^{\pi}(x^{\prime})\}.$ (8.21)
Iteratively updating $F^{\pi}(x)$ using this equation converges to the fixed point, which approaches the true HJ reachability function as $\gamma\to 1$ [fisac2019bridging].
In the following, we present the trajectories and feasible regions for MPC under the pointwise and CBF constraints, as well as the results for RL under the HJ reachability constraint. To visualize the results on the 2D plane, we fix the initial velocity as $v_{0}=1\mathrm{m/s}$ and the initial heading angle as $\theta_{0}=\pi/2$. Figure 8.18 shows the state trajectories and feasible regions of MPC under pointwise constraints with different horizons. Similar to the emergency braking task, short-sighted pointwise constraints lead to smaller EFRs, with the trajectories colliding with the obstacle. As the horizon increases, more trajectories become safe, the EFR expands, and the IFR approaches the EFR.
Figure 8.18: State trajectories and feasible regions of MPC under pointwise constraints.
Figure 8.19 shows the state trajectories and feasible regions of MPC under CBF constraints with different parameters. As the parameter $k$ decreases, the EFR expands. It is worth noting that the IFR always equals the EFR under CBF constraints, consistent with our analysis in Section 6.2.1. An interesting phenomenon is that although the initial states of some trajectories are infeasible, these trajectories can still avoid collision with the obstacle. This is because the virtual CBF constraint is more conservative than the real-time constraint, and violating the CBF constraint does not necessarily lead to collision in the real-time domain. This conservativeness can also be observed from the smaller EFR under CBF constraints compared to pointwise constraints with long horizons in Figure 8.18.
Figure 8.19: State trajectories and feasible regions of RL under HJR constraints.
Figure 8.20 shows the state trajectories and feasible regions of RL under the HJ reachability constraint at convergence. The EFR and IFR are almost identical, consistent with our analysis in Section 6.3.2. The trajectories successfully avoid collision with the obstacle and the EFR is significantly larger than that under CBF constraints, close to that of the pointwise constraint with a long horizon. This demonstrates the effectiveness of the learning-based method for approximating the optimal feasibility function. It also shows the advantage of CA-type learning-based feasibility functions: they can achieve larger feasible regions with less conservativeness compared to CIS-type analytical feasibility functions while avoiding the computational complexity of long-horizon pointwise constraints.
Figure 8.20: State trajectories and feasible regions of MPC under CBF constraints.
Chapter 9 Conclusion
@@ 8.2 Unicycle obstacle avoidance
This paper proposes a feasibility theory that applies to both MPC and RL by decoupling states, constraints, and policies in constrained OCPs. We reveal that feasibility depends on the combination of these three elements, extending the traditional viewpoint of OCP-specific feasibility theories in MPC. Based on separated definitions of initial and endless, state and policy feasibility, we analyze the containment relationships between different feasible regions, which enables us to describe feasibility under arbitrary combinations of states, constraints, and policies. After that, we provide virtual-time constraint design rules along with a practical design tool called feasibility function that helps achieve the maximum feasible region. The feasibility function is categorized into control invariant set and constraint aggregation, which lead to different kinds of virtual-time constraints. We review most existing constraint formulations and point out that they are essentially applications of feasibility functions in different forms. Finally, we demonstrate our feasibility theory in an emergency braking control task by visualizing initially and endlessly feasible regions of MPC and RL policies under different kinds of virtual-time constraints.
Acknowledgements.
Yujie Yang, Zhilong Zheng, and Shengbo Eben Li are supported by the National Natural Science Foundation of China (Grant No. 92582205) and the Beijing Natural Science Foundation (Grant No. L257002).
[…]